From the data
x 1 1.5 2 2.5 3
f(x) 1 .5 1.3 .8 1.5
interpolate the value of f(2.8) using the Newton’s backward difference formula.
1
Expert's answer
2020-04-28T16:31:58-0400

Newton’s backward difference table


x 1 1.5 2 2.5 3

f(x) 1 0.5 1.3 0.8 1.5

"\u2207" -0.5 0.8 -0.5 0.7

"\u2207^{2}" 1.3 -1.3 1.2

"\u2207^{3}" -2.6 2.5

"\u2207^{4}" 5.1


"f(x)=y_{n}\u200b(x)+p\u2207y_{n}\u200b(x)+\\frac{p(p+1)\u2207^{2}y_{n}\u200b(x)\u200b}{2!}+\\frac{p(p+1)(p+2)\u2207^{3}y_{n}\u200b(x)}{3!\u200b}+\\frac{p(p+1)(p+2)(p+3)\u22074yn\u200b(x)}{4!}\u200b....\\\\where\\phantom{i}p=\\frac{x\u2212x_{0}}{h\u200b\u200b}\\\\p=\\frac{2.8\u22123}{0.5}\\\\\u200bp=\u22120.4,\\phantom{i}y_{n}\u200b=1.5;\\phantom{i}\u2207y_{n}\u200b(x)=0.7;\\phantom{i}\u2207^{2}y_{n}\u200b(x)=1.2;\\phantom{i}\u2207^{3}y_{n}\u200b(x)=2.5;\u2207^{4}y_{n}\u200b(x)=5.1\\\\f(2.8)=1.5+(\u22120.4)0.7+\\frac{(\u22120.4)(\u22120.4+1)1.2\u200b}{2!}+\\frac{(\u22120.4)(\u22120.4+1)(\u22120.4+2)2.5\u200b}{3!}+\\frac{(\u22120.4)(\u22120.4+1)(\u22120.4+2)(\u22120.4+3)5.1\u200b}{4!}\\\\f(2.8)=1.5\u22120.28\u22120.144\u22120.16\u22120.21216\\\\f(2.8)=0.70384"